AMPLITUDE, PERIOD, PHASE SHIFT AND VERTICAL SHIFT (TRIGONOMETRY)
B. Determine the amplitude, period, phase shift, and vertical shift for each equation.
1. y = 1/4 sin2x
2. y = cos2x - 5
(Show full solution)
Thank You! :)

Answer by greenestamps(13334)



Put each equation in standard form:
y=(a)(sin(b(x-c)))+d
amplitude: a period: (2pi)/b phase shift (horizontal shift): c vertical shift: d
1. y=(1/4)sin(2x) --> y=(1/4)(sin(2(x-0)))+0 a=1/4; b=2; c=0; d=0
or y=(a)(cos(b(x-c)))+d
2. y=cos(2x)-5 --> y=(1)(cos(2(x-0)))+(-5) a=1; b=2; c=0; d=-5